OpenStudy (anonymous):

State Hooke's law. A spring will stretch 12 cm when a 250-g mass is hung on it. If two springs are hung side by side and both support the 250-g mass, how much will each spring stretch? Explain.

OpenStudy (anonymous):

When two springs are side by side like you describe , each spring will stretch the same amount, x, as given by the equation: |Fs| = k1*x+k2*x, or |Fs| = (k1+k2)*x The mass is motionless, so |Fs| = mg mg = (k1+k2)*x x = mg/(k1+k2) The length of stretch depends on the spring constant of each spring.

OpenStudy (anonymous):

How do I find out the spring constant?

OpenStudy (anonymous):

Since "A spring will stretch 12 cm when a 250-g mass is hung on it," Fs = kx --> mg = kx --> k = mg/x --> k=(.250 kg)(9.8 N/kg )/(12cm) = .204 N/cm I guess the question wants to assume the two springs are equal, so x = mg/(2k) --> x (two springs) = .250kg*9.8N/kg/[2*.204 N/cm] = 6 cm The elongation length halves.

OpenStudy (anonymous):

I honestly am not comprehending what I'm seeing. Not at all.

OpenStudy (anonymous):

Force exerted by the spring = Fs k = spring constant x = length of stretch Fs=kx rearrange to put the focus on length of stretch: x = Fs/k In our case, Fs = weight of the "mass", and that stays the same. So the question is basically asking -- if Fs is constant, how does doubling k, the spring constant, affect the length of stretch? if k doubles and Fs remains the same... x = Fs/(2k) Comparing old x and new x: xbefore:xafter Fs/k:Fs/(2k) 1:1/2 Therefore the same weight stretches the spring half as much when the weight is suspended on two springs side by side.

OpenStudy (anonymous):

I'm giving up... /: thank you though

OpenStudy (jtfan2000):

Simple 2 springs at 12 cm is 2x12 = 24 cm